# references.bib A Learning-based Quadcopter Controller with Extreme Adaptation

### v1

@@ III Implementation Details
Simulation Environment. We use the Flightmare simulator [song2020flightmare] to train and test our control policies. We implement the same high-level controller in [mueller2018multicopter] to generate high-level commands at the level of body rates and mass-normalized collective thrust for our low-level controller to track. It is designed as a cascaded linear acceleration controller with desired acceleration mimicking a spring-mass-damper system with natural frequency 2rad/s and damping ratio 0.7. The desired acceleration is then converted to the desired total thrust and the desired thrust direction, and the body rates are computed from this as proportional to the attitude error angle, with a time constant of 0.2s. The high-level controller’s inputs are the platform’s state (position, rotation, angular, and linear velocities) and the reference position, velocity, and acceleration from the generated trajectory at the simulated time point. The policy outputs individual motor speed commands, and we model the motors’ response using a first-order system with a time constant of 10ms. Each RL episode lasts for a maximum of 5s of simulated time, with early termination if the quadcopter height drops below 2cm (equivalent to a crash to the ground), or the quadcopter’s body rate exceeds 10rad/s. The control frequency of the policy is Hz, and the simulation step is 2ms. We additionally implement an observation latency of 5ms.
@@
Hardware Details. For all of our real-world experiments, we use two quadcopters, which differ in mass by a factor of 3.68, and in arm length by a factor of 3.1. The first one, which we name large quadrotor has a mass of 985g, a size of 17.7cm in arm length, a thrust-to-weight ratio of 3.62, a diagonal inertia matrix of [0.004, 0.008, 0.012]kgm2 (as expressed in the z-up body-fixed frame), and a maximum motor speed of 1000rad/s. The second one, small quadrotor, has a mass of 267g, a size of 5.8cm in arm length, a thrust-to-weight ratio of 3.23, a diagonal inertia matrix of [259e-6, 228e-6, 285e-6]kgm2, and a maximum motor speed of 6994rad/s. For each of our platforms, we use a Qualcomm Robotics RB5 platform as the onboard computer which runs the high-level control and our deployed policy, and a mRo PixRacer as the flight control unit. We use as high-level a PID controller which takes as input the goal position, velocity, and acceleration and outputs the mass normalized collective trust and the body rates. An onboard Inertia Measurement Unit (IMU) measures the angular velocity and the acceleration of the robot, which is low-pass filtered to reduce noise and remove outliers. The high-level commands of the collective thrust and the body rates, and the low-level measurement of the angular rates and the acceleration are fed into the deployed policy as inputs. The policy outputs motor speed commands, which are sent to the PixRacer via the UART serial port and subsequently tracked by off-the-shelf electronic speed controllers.
@@
Network Architecture and Training Procedure. The base policy is a 3-layer MLP with 256-dim hidden layers. This takes the drone state and the vector of intrinsics as input to produce motor speeds. The environment factor encoder is a 2-layer MLP with 128-dim hidden layers. The policy and the value function share the same factor encoding layer. The adaptation module projects the latest 100 state-action pairs into a 128-dim representation, with the state-action history initialized with zeros. Then, a 3-layer 1-D CNN convolves the representation across time to capture its temporal correlation. The input channel number, output channel number, kernel size and stride of each CNN layer are [32, 32, 8, 4], [32, 32, 5, 1], [32, 32, 5, 1]. The flattened CNN output is linearly projected to estimate . For RL, we train the base policy and the environment encoder using PPO [schulman2017proximal] for 100M steps in PyTorch. We use the reward described in Section II-B3. Policy training takes approximately 1.5 hours on an ordinary desktop machine with 1 GPU. We then train the adaptation module with supervised learning by rolling out the student policy. We train with the ADAM optimizer to minimize MSE loss. We run the optimization process for 10M steps, training on data collected over the last 1M steps. Training the adaptation module takes approximately 20 minutes. Both networks are trained with the deep learning framework PyTorch. For more efficient inference and resource allocation on the onboard computer, we use Mobile Neural Network (MNN) [jiang2020mnn, proc:osdi22:walle] to convert trained models to MNN formats to optimize their inference speed and overhead.
@@ IV-A Baselines Setup
We compare our approach with a set of baselines in the simulation. The task is to evaluate the tracking performance of a randomly sampled quadcopter along random trajectories. We randomize quadcopters according to our design-informed domain randomization technique outlined in Section II-B4. The testing range is listed in Table II and a sample of typical desired trajectories is shown in Figure 5. We choose a nominal quadcopter model , which is obtained by setting when sampling without noise added.
We choose several different sets of high-level and low-level controllers as baselines from prior work. We choose PID-PD∗, with PID as the high-level controller and a low-level proportional-derivative controller with access to the ground truth model parameters of the sampled vehicle (PD∗); and PID-PDn with the low-level proportional-derivative controller only using parameters of to control all the sampled quadcopters (PDn). We choose these two baselines in particular because the PID-PD∗ can serve as the expected performance upper bound since it has access to ground-truth model parameters, and the PID-PDn can serve as the lower bound since it uses only the nominal model aiming to control all sampled vehicles. The two baselines can be used for the sanity check for our other proposed baselines and our approach, since their performance should be within the range bounded by these two methods.
@@
In addition, we choose -PDn with as a model-based adaptive high-level controller ([cao2008design, hovakimyan2010l1, hanover2021performance]) and PDn as the low-level controller; and PID- with as the low-level adaptive controller. We design our method as a low-level controller under the assumption that the adaptation across the diverse parametric range as Table II is more effective at this control level. In other words, an adaptive high-level controller solely cannot sufficiently account for the model disparity in our problem, which aims to control quadcopters with significant difference in design and actuators. The purpose of baselines is to validate our assumption by comparing their performance when the adaptation is at a different control level.
@@
At the beginning of each experiment, the quadcopter is spawned with a hovering state. The trajectory to track is generated with the motion primitive generation algorithm [mueller2015trajectory] with the end condition sampled from the test range in Table II. The experiment is considered successful if the position tracking error is within 2m at every point of the trajectory. This failure condition provides a clear distinction between a successful flight and a complete failure.
@@ IV-B Results
Success Rate | Position | Velocity | |||
(m) | (m/s) | ||||
PID-PDn | 22% | 0.5100.372 | 0.8451.066 | ||
-PDn | 62% | 0.1860.167 | 0.2780.392 | ||
PID- | 77% | 0.2210.242 | 0.3570.481 | ||
PID-Ours | 100% | 0.1540.079 | 0.1170.068 | ||
100% | 0.0610.057 | 0.0590.050 |
The results of the simulation experiments are reported in Table III. We compare the five approaches under three metrics: (i) the success rate, (ii) the root-mean-square error (RMSE) in position tracking, and (iii) the RMSE in velocity tracking. We rank the methods according to the success rate and the tracking performance.
Given the very large amount of quadcopter variations, PID-PDn with only access to the nominal model achieved the lowest success rate and the largest tracking error. In contrast, PID-PD∗ has a 100% success rate with the lowest tracking error, since it uses the ground-truth parameters of the quadcopter in computing the control inputs. Without access to the ground truth parameters as PID-PD∗ but with adaptation to the unknown dynamics, the flight performance of controllers significantly increases. However, with adaptation at high-level, the -PDn achieves a lower success rate than its counterpart PID- with adaptation at low-level. Since tracking errors are only computed in successful runs, the -PDn achieves a slightly lower tracking error. This result has shown that an adaptive low-level controller tends to perform better with the large model disparity across the platforms, which justifies our assumption for our controller design.
PID- is the strongest baseline after PID-PD in our simulations. However, its success depends on knowing the reference model , while our method operates without this prior knowledge. Despite this, our method achieves a 100% success rate like PID-PD, with only a slightly higher tracking error. This difference is expected, as PID-PD uses the ground-truth model parameters, whereas ours does not.
@@ IV-C Generalization
The success rate and the position and velocity tracking error distribution are reported in Figure 7. Our method achieve above 95% success rate until where it drops to 84%. In contrast, all other baselines, except for PID-PD∗, exhibit significant performance degradation as the model mismatch from the nominal model increases. In addition, the average position tracking error at of our method is still close to that at , with only 4.6% increase. Compared to the strongest baseline PID- in Section IV-B whose position tracking error has grown by 481.1% compared to that at the nominal model.
@@ V Hardware Experiments
In this section, we transition from simulation to hardware experiments. We first investigate the sim-to-real correlation to ensure the validity of our simulation results. Subsequently, we conduct a comparative analysis on disturbance rejection tasks, comparing our method against the best baseline identified in our simulation tests. This section aims to validate the effectiveness and robustness of our approach in real-world conditions.
@@ V-A Sim-to-Real Correlation
We validate our simulation results through two sets of hardware and simulation experiments to examine the simulation-to-reality gap. We fly the large quadrotor along circular trajectories with our control framework PID-Ours at 3 speed settings: Slow, Medium and Fast. Subsequently, we simulate the same flight paths of the same vehicle with our method again using the Flightmare simulator [song2020flightmare]. The trajectories involve circling with a 1-meter radius with completion in different durations: 8s (Slow), 4s (Medium), and 3s (Fast). The results, illustrated in Figure 8, show the distribution of position and velocity tracking errors for both sets of experiments across all speed settings. We also compute the Pearson Correlation Coefficient [freedman2007statistics, benesty2009pearson], which shows a moderate positive correlation (0.652) between the position tracking errors in the simulation and real-world tests, with a statistically significant P-value of 0.002. These findings suggest that our simulator captures the dynamics of the real world reasonably well, supporting the reliability of our simulation results.
@@ V-B Baseline Comparison
Vehicle | Low-level | Thrust | Average Angular Vel. | |
Controller | RMSE (m/s2) | RMSE (rad/s) | ||
Disturb Free | small | PD∗ | 0.132 | 0.339 |
Ours | 0.280 | 0.721 | ||
large | PD∗ | 2.327 | 0.296 | |
Ours | 3.325 | 0.413 | ||
Wind | small | PD∗ | 2.477 | 1.365 |
Ours | 1.659 | 0.429 | ||
large | PD∗ | 2.916 | 0.543 | |
Ours | 3.549 | 0.523 | ||
Off-center Payload | small | PD∗ | 0.730 | 1.360 |
Ours | 0.679 | 0.510 | ||
large | PD∗ | 3.420 | 0.697 | |
Ours | 2.808 | 0.456 | ||
Thrust Loss | small | PD∗ | 0.689 | 1.408 |
Ours | 0.403 | 0.576 | ||
large | PD∗ | Fail | Fail | |
Ours | 2.167 | 0.355 |
@@
The high-level controller for both methods is a PID controller, the same as in Section III. The only variation in the two control frameworks is the low-level controller. Therefore, we compare the two approaches by their high-level command tracking performance. The comparison is conducted under two metrics: the average tracking error of the (i) mass normalized thrust and (ii) angular velocity given the high-level controller’s commands. We define a failure as a situation where the human operator has to intervene to prevent the quadcopter from crashing. The results of these experiments are reported in table IV. Note that in nearly all experiments, the thrust tracking RMSE for large quadrotor is much higher than that for the small quadrotor. This discrepancy arises because the large quadrotor has more powerful actuators, which can induce greater vibrations in the system, subsequently affecting the accelerometer readings. Therefore, this difference does not necessarily indicate that the thrust tracking for the large quadrotor is worse than for the small quadrotor. It is more appropriate to compare the performance of different methods within each platform, rather than across them.
@@
Our approach and the PD∗ baseline perform similarly in disturbance-free experiments, with the PD∗ controller slightly outperforming ours. The latter difference in performance is justified since the PD∗ controller is specifically tuned for each quadcopter. Our method significantly outperforms the model-based PD∗ in both metrics in the presence of off-center payload and thrust loss. In particular, PID-PD∗ experiences a total failure in the case of thrust loss on the large quadrotor platform. Both disturbances create a large model mismatch from the nominal model that the PID-PD∗ uses. Our method is able to adapt to the mismatch well with a similar high-level command tracking error compared to that at the disturbance-free case. Conversely, the PID-PD∗ controller is not as adaptive. Indeed, its tracking error is up to 4.53 times higher for thrust tracking and up to 3.15 times for angular velocity tracking. Finally, the purpose of wind experiments is to evaluate our controller’s performance under non-constant disturbances. Note that such time-varying disturbances were not present during training. Our controller is robust to wind disturbances with a comparable tracking performance as the PID-PD∗ on the large quadrotor and a significantly smaller tracking error on the small quadrotor. The small size and weight of this platform make it more susceptible to the interaction of the wind with its body and rotors, which can alter its aerodynamic properties, such as affecting the effective angle of attack on the rotors. Therefore, the model mismatch in terms of alteration in aerodynamic properties is more prominent on the small quadrotor than on the large quadrotor. Our controller can adapt well to such disturbances.
VI Conclusion

### v2

@@ III Implementation Details
Simulation Environment. We use the Flightmare simulator [37] to train and test our control policies. We implement the same high-level controller in [38] to generate high-level commands at the level of body rates and mass-normalized collective thrust for our low-level controller to track. It is designed as a cascaded linear acceleration controller with desired acceleration mimicking a spring-mass-damper system with natural frequency 2rad/s and damping ratio 0.7. The desired acceleration is then converted to the desired total thrust and the desired thrust direction, and the body rates are computed from this as proportional to the attitude error angle, with a time constant of 0.2s. The high-level controller’s inputs are the platform’s state (position, rotation, angular, and linear velocities) and the reference position, velocity, and acceleration from the generated trajectory at the simulated time point. The policy outputs individual motor speed commands, and we model the motors’ response using a first-order system with a time constant of 10ms. Each RL episode lasts for a maximum of 5s of simulated time, with early termination if the vehicle loses more than 10m from its starting height, or the quadcopter’s body rate exceeds 10rad/s. The control frequency is 500Hz, which is also the simulation step. We additionally implement an measurement latency of 5ms.
@@
Hardware Details. For all of our real-world experiments, we use two quadcopters, which differ in mass by a factor of 3.68, and in arm length by a factor of 3.1. The first one, which we name large quadrotor has a mass of 985g, a size of 17.7cm in arm length, a thrust-to-weight ratio of 3.62, a diagonal inertia matrix of [0.004, 0.008, 0.012]kgm2 (as expressed in the z-up body-fixed frame), and a maximum motor speed of 1000rad/s. The second one, small quadrotor, has a mass of 267g, a size of 5.8cm in arm length, a thrust-to-weight ratio of 3.23, a diagonal inertia matrix of [259e-6, 228e-6, 285e-6]kgm2, and a maximum motor speed of 6994rad/s. For each of our platforms, we use a Qualcomm Robotics RB5 platform as the onboard computer which runs the high-level control at 50Hz and our deployed policy at 500Hz, and a mRo PixRacer as the flight control unit. We use as high-level a PID controller which takes as input the goal position, velocity, and acceleration and outputs the mass normalized collective trust and the body rates. An onboard Inertia Measurement Unit (IMU) measures the angular velocity and the acceleration of the robot, which is low-pass filtered to reduce noise and remove outliers. The high-level commands of the collective thrust and the body rates, and the low-level measurement of the angular rates and the acceleration are fed into the deployed policy as inputs. The policy outputs motor speed commands, which are sent to the PixRacer via the UART serial port and subsequently tracked by off-the-shelf electronic speed controllers. The real-world experiment is performed indoors with a motion capture system running. The motion capture is only used for evaluating the system’s performance in experiments and for feedback to the high-level controller, but it does not provide any feedback to our policy.
@@
Network Architecture and Training Procedure. The base policy is a 3-layer MLP with 256-dim hidden layers. This takes the drone state and the vector of intrinsics as input to produce motor speeds. The environment factor encoder is a 2-layer MLP with 128-dim hidden layers. The policy and the value function share the same factor encoding layer. The adaptation module projects the latest 100 state-action pairs into a 128-dim representation, with the state-action history initialized with zeros. We selected a window size of 100 as it provides good performance while keeping the network lightweight. Then, a 3-layer 1-D CNN convolves the representation across time to capture its temporal correlation. The input channel number, output channel number, kernel size and stride of each CNN layer are [32, 32, 8, 4], [32, 32, 5, 1], [32, 32, 5, 1]. The flattened CNN output is linearly projected to estimate the intrinsics vector . For RL, we train the base policy and the environment encoder using PPO [39] for 100M steps in PyTorch. We use the reward described in Section II-C. Policy training takes approximately 1.5 hours on an ordinary desktop machine with 1 NVIDIA GeForce RTX 4060 GPU. We then train the adaptation module with supervised learning by rolling out the student policy. We train with the ADAM optimizer to minimize MSE loss. We run the optimization process for 10M steps, training on data collected over the last 1M steps. Training the adaptation module takes approximately 20 minutes. Both networks are trained with the deep learning framework PyTorch. For more efficient inference and resource allocation on the onboard computer, we use Mobile Neural Network (MNN) [40, 41] to convert trained models to MNN formats to optimize their inference speed and overhead. Table III presents the average inference time measured on the RB5 platform for onboard computation. To meet the required 500Hz control frequency, the combined inference time of the controller policy and adaptation module must be less than or equal to 2ms. MNN satisfies this requirement with a total average inference time of 0.165ms, whereas PyTorch exceeds the limit with a total of 78.850ms, making real-time onboard inference infeasible.
PyTorch | MNN | ||
Low-level Controller () | 73.6262.692 | 0.0780.003 | |
Adaptation Module () | 5.2240.708 | 0.0870.005 |
@@ IV-A Baselines Setup
We compare our approach with a set of baselines in the simulation. The task is to evaluate the tracking performance of a randomly sampled quadcopter along random trajectories. We randomize quadcopters according to our design-informed domain randomization technique outlined in Section II-E. The testing range is listed in Table II and a sample of typical desired trajectories is shown in Figure 4. We choose a nominal quadcopter model , which is obtained by setting when sampling without noise added.
We choose several different sets of high-level and low-level controllers as baselines from prior work. We include the implementation details of all baselines in Appendix A. Most baseline names follow the format high-level controller-low- level controller, with exceptions explicitly noted. First, we include PID-PD∗ and PID-PDn as reference baselines. PID-PD∗ serves as an upper performance bound, as its low-level PD controller has access to the ground-truth model parameters of each sampled quadcopter. This makes it equivalent to the expert controller used during training, as described in Section II-D. In contrast, PID-PDn serves as a lower performance bound, since its low-level PD controller only uses nominal parameters from the model and does not adapt to individual quadcopters. These two baselines provide a sanity check: any effective adaptation method should achieve performance within this range. To ensure that the adaptation task is handled entirely by the low-level controller, all high-level controllers in both the baselines and our framework use the same control parameters across all tested vehicles. These control gains are tuned specifically to optimize performance on the nominal model .
@@
Next, we include -PDn and PID- as additional baselines for evaluating our method. -PDn employs the adaptive high-level controller ([4, 5]) to compensate for model uncertainties while keeping the low-level PD controller fixed. In contrast, PID- applies as a low-level adaptive controller. These baselines serve a dual purpose: they act as additional benchmarks for evaluating our approach while also validating our core assumption about adaptation. We design our method as a low-level controller under the assumption that adaptation across the diverse parametric range in Table II is more effective when handled at this control level. In other words, an adaptive high-level controller alone cannot sufficiently compensate for the model disparity in our problem, which involves controlling quadcopters with significant differences in design and actuators. The baselines allow us to test this assumption by comparing their performance when adaptation occurs at different control levels.
Finally, we compare our method to state-of-the-art adaptive controllers. We include Geo-A, a geometric adaptive controller that operates at both high and low levels [12] so is an exception to our naming convention. We also include PID-INDI-A, which combines a PID high-level controller with a low-level adaptive INDI controller [13]. These baselines provide further context on how our method compares to established nonlinear adaptive control techniques.
@@
At the beginning of each experiment, the quadcopter is spawned with a hovering state. The trajectory to track is generated with the motion primitive generation algorithm [33] with the end condition sampled from the test range in Table II. The experiment is considered successful if the position tracking error is within 2m at every point of the trajectory.
@@ IV-B Results
Success | Max Pos. | Position | Velocity | |||
Rate | Err (m) | (m) | (m/s) | |||
PID-PDn | 22% | 1.565 | 0.5100.372 | 0.8451.066 | ||
-PDn | 62% | 1.105 | 0.1860.167 | 0.2780.392 | ||
Geo-A | 64% | 1.033 | 0.1280.160 | 0.1740.263 | ||
PID-INDI-A | 67% | 0.333 | 0.0630.060 | 0.0750.082 | ||
PID- | 77% | 1.304 | 0.2210.242 | 0.3570.481 | ||
PID-Ours | 100% | 0.311 | 0.1480.075 | 0.1290.066 | ||
100% | 0.221 | 0.0610.057 | 0.0590.050 |
The results of the simulation experiments are reported in Table IV. We compare the five approaches under three metrics: (i) the success rate, (ii) the maximum position tracking error, (iii) the root-mean-square error (RMSE) in position tracking, and (iv) the RMSE in velocity tracking. We rank the methods according to the success rate and the tracking performance.
Given the very large amount of quadcopter variations, PID-PDn with only access to the nominal model achieved the lowest success rate and the largest tracking error. In contrast, PID-PD∗ has a 100% success rate with the lowest tracking error, since it uses the ground-truth parameters of the quadcopter in computing the control inputs. Without access to the ground truth parameters as PID-PD∗ but with adaptation to the unknown dynamics, the flight performance of controllers significantly increase. However, with adaptation at high-level, the -PDn achieves a lower success rate than its counterpart PID- with adaptation at low-level. Since tracking errors are only computed in successful runs, the -PDn achieves a slightly lower tracking error. This result has shown that an adaptive low-level controller tends to perform better with the large model disparity across the platforms, which justifies our assumption for our controller design.
The two nonlinear adaptive baselines Geo-A and PID-INDI-A both show better tracking performance than controllers. Specifically, PID-INDI-A achieves the lowest tracking error in successful flights among all methods. However, both baselines achieve a similar success rate of approximately 65%. Specifically, we often observe them failing near the boundaries of the adaptation range, where the dynamics deviate significantly from the nominal model. Moreover, their performance relies on prior knowledge of the reference model , and INDI-A in particular also requires access to current motor speed measurements and accurate estimates of angular acceleration and torque, whereas our method operates without such information. Despite this, our method achieves a 100% success rate and the lowest maximum position error among all baselines, with only a slightly higher average tracking error than the expert controller with access to the true parameters.
@@ IV-C Generalization
The success rate and the position and velocity tracking error distribution are reported in Figure 6. Our method achieves near 100% success rate until where it drops to 95%. In contrast, all other baselines, except for PID-PD∗, exhibit significant performance degradation as the model mismatch from the nominal model increases. In addition, the average position tracking error at of our method is still close to that at , with only 37.0% increase. Compared to the strongest baseline PID- in terms of success rate in Section IV-B whose position tracking error has grown by 483.7% compared to that at the nominal model.
@@ V Hardware Experiments
In this section, we transition from simulation to hardware experiments. We first investigate the sim-to-real correlation to ensure the validity of our simulation results. Subsequently, we conduct a comparative analysis on disturbance rejection tasks, comparing our method against the best baseline identified in our simulation tests.
@@ V-A Sim-to-Real Correlation
We validate our simulation results through two sets of hardware and simulation experiments to examine the simulation-to-reality gap. We fly the large quadrotor along circular trajectories with our control framework PID-Ours at 3 speed settings: Slow, Medium and Fast. Subsequently, we simulate the same flight paths of the same vehicle with our method again using the Flightmare simulator [37]. The trajectories involve circling with a 1-meter radius with completion in different durations: 8s (Slow), 4s (Medium), and 3s (Fast). The results, illustrated in Figure 7, show the distribution of position and velocity tracking errors for both sets of experiments across all speed settings. We also compute the Pearson Correlation Coefficient [42, 43], which shows a moderate positive correlation (0.652) between the position tracking errors in the simulation and real-world tests, with a statistically significant P-value of 0.002. These findings suggest that our simulator captures the dynamics of the real world reasonably well, supporting the reliability of our simulation results.
@@ V-B Baseline Comparison
Vehicle | Low-level | Thrust | Average Angular | Max Pos. | Position | Velocity | |
Controller | RMSE (m/s2) | RMSE (rad/s) | Err (m) | RMSE (m) | RMSE (m/s) | ||
Disturb Free | small | PD∗ | 0.132 | 0.339 | 0.094 | 0.017 | 0.022 |
Ours | 0.280 | 0.721 | 0.098 | 0.033 | 0.047 | ||
large | PD∗ | 2.327 | 0.296 | 0.108 | 0.046 | 0.132 | |
Ours | 3.325 | 0.413 | 0.122 | 0.057 | 0.197 | ||
Wind | small | PD∗ | 2.477 | 1.365 | 0.245 | 0.175 | 0.328 |
Ours | 1.659 | 0.429 | 0.191 | 0.130 | 0.205 | ||
large | PD∗ | 2.916 | 0.543 | 0.576 | 0.072 | 0.188 | |
Ours | 3.549 | 0.523 | 0.341 | 0.076 | 0.131 | ||
Off-center Payload | small | PD∗ | 0.730 | 1.360 | 0.435 | 0.260 | 0.323 |
Ours | 0.679 | 0.510 | 0.185 | 0.130 | 0.158 | ||
large | PD∗ | 3.420 | 0.697 | 0.703 | 0.404 | 0.564 | |
Ours | 2.808 | 0.456 | 0.253 | 0.167 | 0.215 | ||
Thrust Loss | small | PD∗ | 0.689 | 1.408 | N.A. | 0.470 | 0.389 |
Ours | 0.403 | 0.576 | N.A. | 0.325 | 0.196 | ||
large | PD∗ | Fail | Fail | Fail | Fail | Fail | |
Ours | 2.167 | 0.355 | N.A. | 0.572 | 0.224 |
@@
The high-level controller for both methods is a PID controller, the same as in Section III. The only variation in the two control frameworks is the low-level controller. We asses performance using the same metrics of the simulation experiments. However, to test the controller’s ability to track desired commands, we additionally measure the average tracking error of (i) mass-normalized thrust and (ii) angular velocity based on the high-level controller’s commands. We define a failure as a situation where the human operator has to intervene to prevent the quadcopter from crashing. The results of these experiments are reported in table V. Note that in nearly all experiments, the thrust tracking RMSE for large quadrotor is much higher than that for the small quadrotor. This discrepancy arises because the large quadrotor has more powerful actuators, which can induce greater vibrations in the system, subsequently affecting the accelerometer readings. Therefore, this difference does not necessarily indicate that the thrust tracking for the large quadrotor is worse than for the small quadrotor. It is more appropriate to compare the performance of different methods within each platform, rather than across them.
@@
Our approach and the PD∗ baseline perform similarly in disturbance-free experiments, with the PD∗ controller slightly outperforming ours. The latter difference in performance is justified since the PD∗ controller is specifically tuned for each quadcopter. Our method significantly outperforms the model-based PD∗ in both metrics in the presence of off-center payload and thrust loss. In particular, PID-PD∗ experiences a total failure in the case of thrust loss on the large quadrotor platform. Both disturbances create a large model mismatch from the nominal model that the PID-PD∗ uses. Our method is able to adapt to the mismatch well with a similar high-level command tracking error compared to that at the disturbance-free case. Conversely, the PID-PD∗ controller is not as adaptive. Indeed, its tracking error is up to 4.53 times higher for thrust tracking and up to 3.15 times for angular velocity tracking. Finally, the purpose of wind experiments is to evaluate our controller’s performance under non-constant disturbances. Note that such time-varying disturbances were not present during training. Our controller is robust to wind disturbances with a comparable tracking performance as the PID-PD∗ on the large quadrotor and a significantly smaller tracking error on the small quadrotor. The small size and weight of this platform make it more susceptible to the interaction of the wind with its body and rotors, which can alter its aerodynamic properties, such as affecting the effective angle of attack on the rotors. Therefore, the model mismatch in terms of alteration in aerodynamic properties is more prominent on the small quadrotor than on the large quadrotor. Our controller can adapt well to such disturbances. Across all sets of experiments, position and velocity errors correlate with high-level tracking command errors, as larger high-level errors lead to compounded position and velocity inaccuracies. This further emphasizes the importance of robustness and adaptivity of the low-level controller.
V-C Adaptation Analysis
We consider the controller performance for an off-center payload, using the large quadcopter. The result is shown in Figure 9. During takeoff, all components of exhibit variations as the quadrotor stabilizes. The vehicle transitions from stationary on the ground to takeoff at the commanded velocity within approximately 0.2 seconds, aligning with the adaptation module’s time window (100 state-action pairs at 500Hz). This adaptation process is significantly faster than approaches that rely on online system identification [44, 45], which typically require 10–15 seconds of flight to converge to an accurate model estimate.
When the payload is added, multiple components undergo a sharp transient response, indicating an update in the controller’s internal representation. The highlighted components show a sustained shift, suggesting that these dimensions capture key physical properties such as mass distribution and inertial variations. In contrast, the average value of and remains largely unchanged after convergence, suggesting that these dimensions encode properties that are unaffected by an off-center payload, such as lateral forces or yaw torque.
VI Ablation Study
Training a low-level controller that adapts across different quadrotors involves various complexities. This section presents an ablation study analyzing the training curves to evaluate the impact of the IL component and reward design. Additionally, we examine how the RL and IL components, and our control-level selection, influence performance.
VI-A Training Curve Analysis
VI-A1 Integration of IL
Mean Episode Reward | Mean Episode Length | |
95% CI | 95% CI | |
IL+RL | -31.7222.02 | 0.970.05 |
RL | -102.5576.39 | 0.790.20 |
We adopt a dual strategy which combines IL and RL and adaptively adjust the relative weight between the two losses. This learning approach improves the training of the low-level controller compared to training solely with RL, as in our previous conference paper [29]. Figure 10 presents an ablation of the IL component of the loss. We also report the average episode reward and length with 95% confidence interval (CI) over the last 10% of training process in Table VI. The policy trained with both IL and RL exhibits steady improvement in both reward and episode length throughout the learning phase, maintaining an episode length close to 1 after approximately 10M steps. In contrast, the RL-only baseline shows less consistent performance with a larger variance. Although its episode length approaches 1 around 20M steps, it soon begins to diverge, indicating instability. Additionally, the RL-only baseline maintains a consistently lower average reward compared to the IL-and-RL method, highlighting inferior tracking performance. As training progresses, the instability worsens, leading to a decline in reward. To further analyze these components, we conduct an ablation study in the context of our simulation experiments and evaluate the roles of the RL and IL components in Section VI-B.
VI-A2 Reward Design
Mean Episode Reward | |
95% CI | |
Torque Tracking | -3.914.67 |
Angular Velocity Tracking | -8.74 8.15 |
We reward torque tracking instead of angular velocity tracking, as described in Section II-C. While angular velocity is a high-level command and a key observation, our experiments show that torque tracking leads to better sim-to-real transfer.
To support this, we compare policies trained with torque tracking and angular velocity tracking in Figure 11, keeping all other reward components and hyperparameters identical. Table VII reports the average reward over the last 10% of training. The torque-tracking policy exhibits stable improvement throughout training, whereas the angular velocity-tracking policy achieves lower rewards and shows greater fluctuations. Such behavior not only increases oscillation penalties but also degrades sim-to-real transfer, as saturated commands can overheat motors and cause hardware damage. Torque tracking mitigates these issues by providing a more immediate and direct learning signal. Since torque responds directly to commanded motor speed, it enables faster corrections, which is particularly important for low-level controllers operating at 500Hz. By contrast, angular velocity is derived through integration, making its feedback less immediate for RL.
VI-B Simulation Performance
In this section, we present an ablation study to investigate the contribution of individual components in our method design within the context of our simulation experiments. Specifically, we compare our method against three variations: (i) RL-Only, which excludes the IL component and relies entirely on reinforcement learning for training, (ii) IL-Only, which trains the controller solely through imitation learning using DAgger [46] without separation of networks and phases as in our framework, and (iii) End-to-End, which replaces our hierarchical control structure with a unified policy that directly fuses high-level and low-level control, trained using our method. The first two variants provide insight into the significance of individual training components, extending the analysis from the training curves in Section VI-A. The third variation focuses on justifying our design choice for separate control levels. We perform the experiment on the same generalization task described in Section IV-C. The success rate and the distribution of position tracking errors are presented in Figure 12. Note that the results for Ours are the same as shown in Figure 6.
VI-B1 RL and IL Ablation
Both RL-Only and IL-Only experience a decline in success rate as the model mismatch increases beyond the training set, ultimately dropping to around 60%. RL-Only demonstrates a slightly higher success rate than IL-Only. In terms of tracking error, IL-Only achieves the smallest distribution among the methods. However, both approaches exhibit increased tracking error as the model mismatch grows. At , the average tracking error for IL-Only increases by 92.4%, while for RL-Only, it grows by 61.4%, compared to their respective values at . In contrast, our method, which combines RL and IL, maintains a near 100% success rate across all levels of model mismatch while achieving consistent tracking performance. This raises the question: how can the combination of two methods, each with limited generalization capability, yield better results?
We hypothesize that it is because IL and RL optimize distinct objectives. IL focuses on minimizing instantaneous tracking error, while RL targets both tracking error and success rate. RL achieves this dual optimization by incorporating a reward signal for tracking error, as described in Section II-C, and utilizing episode termination conditions to influence success rate. As a result, RL may accept higher tracking errors to extend episode length, thereby improving success rates. This explains why RL-Only outperforms IL-Only in terms of success rate. When combined, RL can prioritize success rate optimization, as the IL loss naturally handles tracking error minimization. This complementary relationship allows the two methods to reinforce each other, resulting in improved overall performance. Future work could explore this hypothesis further, investigating the underlying mechanisms and conditions that enable a synergy between IL and RL.
VI-B2 End-to-End Evaluation
End-to-End is trained using our proposed method but as a unified policy that combines high-level and low-level control. Its input includes the same features as our policy, along with additional information: position (), velocity (), and the difference between the current and goal positions (). Its output, like our policy, is motor speed, which motivates the name End-to-End. Despite being trained with the same method and without ablations, End-to-End demonstrates the lowest success rate and the largest average tracking error among all tested variations. In contrast, the other variations (RL-Only, IL-Only and Ours) all utilize a low-level controller. The failure of End-to-End learning likely comes from the challenge of jointly encoding both high-level trajectory tracking and low-level motor adaptation within a single policy. Maintaining a modular structure allows each level to focus on distinct aspects of control, facilitating generalization beyond the training set.
Moreover, when compared to state-of-the-art model-based methods in Figure 6, its performance is similarly poor. At , End-to-End achieves a success rate of only 28%, which is marginally better than PID-PDn and -PDn, both of which lack adaptation in the low-level controller. This comparison in parallel also highlights the importance of separating control levels and employing an adaptive low-level controller to handle the large model disparities defined in our problem.
VII Conclusion